OpenStudy (anonymous):

I need help with another equation solve x-5/x+2= 4/7

OpenStudy (anonymous):

"Cross-multiply" : 7x - 35 = 4x +8 3x = 43 x = 43/3

OpenStudy (anonymous):

thank u, I really have a hard time with math